OpenStudy (anonymous):

The interior angles of a polygon equal 360°. How many sides does this polygon have?

OpenStudy (anonymous):

A 3 B 4 C 6

OpenStudy (whpalmer4):

The sum of the interior angles of an n-sided polygon is \[(n-2)*180^\circ\]

jimthompson5910 (jim_thompson5910):

Solve for n S = 180(n-2) 360 = 180(n-2)

OpenStudy (whpalmer4):

For a triangle, we have \(n=3\), so the sum of the interior angles of a triangle is \((3-2)*180^\circ = 180^\circ\) and I'm sure you know that to be true, even if you are unaware of the formula.

OpenStudy (anonymous):

then what

OpenStudy (whpalmer4):

Well, as Jim said, solve \[360 = (n-2)*180\]for \(n\)

OpenStudy (whpalmer4):

Or think of what regular polygon has 4 sides instead of 3, and what its angles add up to...